OpenStudy (poohbear):

Your family drives to New York ( a distance of 240 miles ) and uses 15 gallons of gasoline. If you had a car that traveled 20% more miles per gallon, how much less gasoline would you have needed to make this entire trip?

OpenStudy (anonymous):

so if you go 240 miles and you 15 miles per gallon how much gasoline you need is this what you need to know

OpenStudy (anonymous):

240/15=16-20%=12.8

OpenStudy (poohbear):

Isn't it suppose to be adding 20%?

OpenStudy (anonymous):

o then 20%=3.2 + 16 = 19.2

OpenStudy (poohbear):

So, would it be 2 gallons less?

OpenStudy (anonymous):

3.2 more gallons

OpenStudy (poohbear):

I meant the answer.

OpenStudy (anonymous):

15+20% =18 240/18=13.33333333 and you would of need 16 gallons 16-13.33=2.67

OpenStudy (anonymous):

what are the options A B C D

OpenStudy (poohbear):

It's a free response.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

you can compute your gas mileage by division you know you get \(240\div 15=16\)mpg a 20% increase in gas mileage would be \(16\times 1.2=19.2\) so the question comes down to "if you get 19.2 mpg, how many gallons of gas does it take to go 240 miles?"

OpenStudy (anonymous):

divide miles by miles per gallon to get \(240\div 19.2 = 12.5\) gallons,

OpenStudy (anonymous):

so i guess you save two and a half gallons

OpenStudy (poohbear):

Ohhhh, I get it. thanks

OpenStudy (anonymous):

yw